OpenStudy (anonymous):

PLEASE HELP AND EXPLAIN A particle moves along the x-axis so that its velocity at anytime is t greater than or equal to 0 is given by v(t)=1-sin(2pi t) a) find the acceleration a(t) of the particle at any time. b) find all values of t, 0

OpenStudy (anonymous):

a) \(a(t)= v^\prime(t)\) b) find the t when v(t)=0 c) \(x(t)=\int\limits_{0}^{t}v(s)ds=\int\limits_{0}^{t}\left(1-\sin(2\pi s)\right) ds.\) Work out the integral.

OpenStudy (anonymous):

thanks